ok i need a little help, I just need the equation if possible, that is where I am having the most trouble, once I have the equation I can usually take it from there.
You invested $4000 in two accounts. In the first part of the investment you earned 4% simple interest. on the second part of the investment, you lost 3%. Combining earnings and losses, the annual incomes from these investments was $55. How much was invest at each rate?

Answer by stanbon(75887)

You invested $4000 in two accounts. In the first part of the investment you earned 4% simple interest. on the second part of the investment, you lost 3%. Combining earnings and losses, the annual incomes from these investments was $55. How much was invest at each rate?
--------------------------------
With two equations:
x + y = 4000
0.04x-0.03y= 55
If you use two equations you solve for x and y.
---------------------------
With one equation
0.04x - 0.04(4000-x) = 55
If you use one equation you solve for x and 4000-x.
========================================================
Cheers,
Stan H.
============================